# DMS*: Minimizing Makespan for Multi-Agent Combinatorial Path Finding

### v1

@@ IV-B DMS* Algorithm
After generating the successor labels $L_{succ}$ of $l$, for each label $l^{\prime}\in L_{succ}$, DMS* checks for conflicts between agents during the transition from $v(l)$ to $v(l^{\prime})$, and store the subset of agents in conflict in the conflict set $I_{C}(l^{\prime})$. DMS* then invokes a procedure BackProp (Alg. 3) to back propagate $I_{C}(l^{\prime})$ to its ancestor labels recursively so that the conflict set of these ancestor labels are modified, and labels with modified conflict set are re-added to OPEN and will be re-expanded. DMS* maintains a back_set($l$) for each label $l$, which is a set of pointers pointing to the predecessor labels to which the back propagation should be conducted. Intuitively, similarly to [10, 8], the conflict set of labels are dynamically enlarged during planning when agents are detected in conflict. The conflict sets of labels determine the sub-graph within the joint graph $\mathcal{G^{W}}$ that can be reached by DMS*, and DMS* always attempts to limit the search within a sub-graph of $\mathcal{G^{W}}$ as small as possible.
Afterwards, if the conflict set $I_{C}(l^{\prime})$ is non-empty, then label $l^{\prime}$ leads to a conflict and is discarded. Otherwise, $l^{\prime}$ is checked for dominance against any existing labels in $\mathcal{F}(v(l^{\prime}))$ using Def. 3. If $l^{\prime}$ is dominated by or is equal to any existing labels in $\mathcal{F}(v(l^{\prime}))$, $l^{\prime}$ is pruned, since any future joint path from $l^{\prime}$ can be cut and paste to a label $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates $l^{\prime}$ without worsening the cost to reach goals. Furthermore, for each label $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates or is equal to $l^{\prime}$, the procedure DomBackProp(Alg. 4) is invoked so that the conflict set $I_{C}(l^{\prime})$ is back propagated to $l$, and $l$ is added to the back_set of $l^{\prime\prime}$. By doing so, DMS* is able to keep updating the conflict set of the predecessor labels of $l^{\prime}$ after $l^{\prime}$ is pruned. This ensures that, when needed, the predecessor labels of $l^{\prime}$ will also be re-expanded after $l^{\prime}$ is pruned.
@@
If label $l^{\prime}$ is not pruned, a procedure SimpleHeu is invoked for $l^{\prime}$ to quickly compute a heuristic that roughly estimates the cost-to-go. One possible way to implement SimpleHeu is to first copy $\vec{h}(l)$, where $l$ is the parent of $l^{\prime}$, and then reduce each component of the copied vector by one except for the components that are already zero. This heuristic is still an underestimate of the cost-to-go since all agents can move at most one step closer to their goals in each expansion. Then, the $f_{temp}(l^{\prime})$ can be computed and $l^{\prime}$ is added to OPEN with $f_{temp}(l^{\prime})$ as its priority. Other related data structure including $\mathcal{F}(v(l^{\prime}))$, back_set, $parent$ are also updated correspondingly, and the current search iteration ends.
@@ IV-C Deferred Target Sequencing
DMS* introduces two techniques to defer the target sequencing until absolutely needed. As aforementioned, the first one uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic when a label is generated, and invokes TargetSeq only when that label is popped from OPEN for expansion.
We now focus on the second technique in DMS*. In the previous MS* [8], every time when the search encounters a new label $l$ that is off-policy, inside TargetSeq, the procedure SolveMHPP needs to be invoked for $l$ to find a joint sequence and policy from $l$, which makes the overall computation burdensome, especially when a lot of new labels are generated due to the agent-agent conflict.
@@
Different from MS*, DMS* seeks to defer the call of SolveMHPP inside TargetSeq. For a label $l^{\prime}$, DMS* attempts to avoid calling SolveMHPP for $l^{\prime}$ by re-using the joint sequence $\gamma(l)$ of its parent label $l$ ($l$ is the parent of $l^{\prime}$). As presented in Alg. 2, on Lines 5-6, DMS* first attemps to build a policy from $l^{\prime}$ by following the joint sequence of its parent label $\gamma(l)$ and computes the corresponding cost-to-go $\vec{h}(l^{\prime})$. Agents $i$ that are not in the conflict set $I_{C}(l)$ (i.e., $i\notin I_{C}(l)$) are still along their individual paths as specified by the previously computed policy $\phi_{\gamma(l)}$. Agents $i$ that are in the conflict set $I_{C}(l)$ (i.e., $i\in I_{C}(l)$) consider all possible actions as described in Equation (1) and may deviate from the individual paths specified by the previously computed policy $\phi_{\gamma(l)}$. Therefore, DMS* needs to go through a check for these agents $i\in I_{C}(l)$: DMS* first computes the $f$-vector by summing up $\vec{g}(l^{\prime})$ and $\vec{h}(l^{\prime})$. Then, if $f^{i}(l^{\prime})$ of some agent $i\in I_{C}(l)$ is no larger than $f_{max}(l)$, DMS* can avoid calling SolveMHPP, since letting the agents follow $\gamma(l)$ in the future will not worsen the objective value. Otherwise, there exists an agent $i\in I_{C}(l)$, whose corresponding $f^{i}(l^{\prime})$ is greater than $f_{max}(l)$, and in this case, DMS* cannot avoid calling SolveMHPP since there may exists another joint sequence from $l^{\prime}$, which leads to a solution joint path with better (smaller) objective value.
IV-D Properties of DMS*
IV-D1 Completeness
@@
This section discusses the properties of DMS*. An algorithm is complete for a problem if finds a solution for solvable instances, and reports failure in finite time for unsolvable instances.
Theorem 1.
DMS* is complete for MCPF-max.
@@
Let $\mathcal{A}=\{0,1\}^{M}$ denote the set of all possible binary vectors and let $\mathcal{S}=\mathcal{G^{W}}\times\mathcal{A}$ denote the Cartesian of the joint graph $\mathcal{G^{W}}$ and $\mathcal{A}$. $\mathcal{S}$ is a finite space. DMS* uses dominance pruning and does not expand the same label twice during the search. As a result, for a unsolvable instance, DMS* enumerates all possible joint paths that starts from $(v_{o},\vec{0}^{M})$ to any reachable $(v,a)\in\mathcal{S}$ and terminates in finite time. For a solvable instance, DMS* conducts systematic search in $\mathcal{S}$ and returns a solution in finite time.
IV-D2 Solution Optimality
@@
Assumption 2.
Line 7 in Alg. 1 never returned true during the search of DMS*.
Theorem 2.
@@
We now explain if Assumption 2 is violated, why DMS* loses the optimality guarantee and how to fix it. In MCPF-max, all agents are “coupled” in the space of binary vectors $\mathcal{A}$ in a sense that a target visited by one agent does not need to be visited by any other agents. As as result, when one agent changes its target sequence by visiting some target that is previously assigned to another agent, all agents may need to be re-planned in order to ensure solution optimality. It means that, when Line 7 in Alg. 1 returns true for label $l^{\prime}$, procedure SolveMHPP needs to be called for $l^{\prime}$ and SolveMHPP may return a new joint sequence $\gamma^{\prime}$. In this new sequence $\gamma^{\prime}$, it is possible that agents within $I_{C}(l)$ visits targets that are previously assigned to agents that are not in $I_{C}(l)$. However, DMS* does not let agents outside $I_{C}(l)$ to take all possible actions as defined in Equation (1). As a result, an optimal solution joint path may lie outside $\mathcal{G}^{W}_{sub}$ and DMS* thus loses the solution optimality guarantee when Assumption 2 is violated.
@@
IV-D3 Solution Bounded Sub-Optimality
@@
Bounded sub-optimal joint sequences lead to inflated heuristic values for labels in DMS*. As DMS* select labels based on $f$-values as A* does, inflated heuristic values lead to bounded sub-optimal solutions for the problem [31].
@@
Theorem 4.
For a solvable instance, if Line 24 in Alg. 1 is replaced with $\vec{f}\leftarrow\vec{g}+(1+w)\vec{h}$ and Assumption 1 and 2 hold, then DMS* returns a $w$-bounded solution joint path.
V Experimental Results
We implement DMS* in Python, and use Google OR-Tool [^5] https://developers.google.com/optimization/routing/routing_tasks as the mHPP solver, which implements the SolveMHPP procedure that is required by DMS*. Limited by our knowledge on Google OR-Tool, we only consider the following type of assignment constraints, where any agent can visit all targets and goals. We use grid maps from the MAPF benchmark data set [32] and set up a 60-second runtime limit for each test instance, where each test instance contains the starts, targets and goals, which are all from the MAPF benchmark data set [32]. All tests run on a MacBook Pro with a Apple M2 Pro CPU and 16GM RAM. It is known from [10] that, for M*-based algorithms, a slightly inflated heuristic can often help improve the scalability of the approach with respect to the number of agents and we set $w=1.1$ in our tests. We set the number of targets $M=20,40,60,80$ in our tests. We compare DMS*, which includes the two proposed technique to defer target sequencing, against MS*, which is the baseline and does not have these two techniques. In other words, MS* here is a naive adaption of the existing MS* [8] algorithm to solve MCPF-max, by using the aforementioned Google OR-Tool to solve min-max mHPP for target sequencing.
Figure 3: The success rates and runtime of both DMS* and MS* (baseline) with varying number of agents and targets in a random 32x32 map. DMS* achieves higher success rates than MS* while requiring less runtime on average.
V-A Varying Number of Agents
We first fix the map to be Random 32x32, and vary the number of agents $N=5,10,15,20$. We measure the success rates, the average runtime to solution and the average runtime for target sequencing per instance. The averages are taken over all instances, including both succeeded instances and instances where the algorithm times out. As shown in Fig. 3, DMS* achieves higher success rates and lower runtime than the baseline MS* for all $N$s. As $M$ increases from $20$ to $80$, both algorithms require more runtime for target sequencing. As $N$ increases from $5$ to $20$, agents have higher density and are more likely to run into conflict with each other. As a result, both algorithms time out for more and more instances. Additionally, MS* spends almost all of its runtime in target sequencing, which indicates the computational burden caused by the frequent call of mHPP solver. In contrast, for DMS*, when $N=5,10$, DMS* spends most of the runtime in target sequencing, while as $N$ increases to $15,20$, DMS* spends more runtime in path planning.
@@ V-A Varying Number of Agents
V-B Different Maps
We then fixed the number of agents to $N=10$ and test both algorithms in several maps of different sizes, as shown in Fig. 4. DMS* outperforms MS* in terms of success rates due to the alleviated computational burden for target sequencing. Furthermore, larger maps do not lead to lower success rates since larger maps can reduce the density of the agents and make the agents less likely to run into conflicts with each other. The Room 32x32 map is relatively more challenging than the other two maps since there are many narrow corridors which often lead to conflicts between the agents.
V-C Experiments with Mobile Robots
We also performed experiments with Husarion ROSbot 2666https://husarion.com/manuals/rosbot/, which are differential drive robots equipped with an ASUS Tinker Board and use a 2D lidar for localization. We used two such robots to perform tests in two different environments, as shown in Fig. 5 (a) and (b) having 2 and 3 targets, respectively. This test verifies that the paths planned by DMS* are executable on real robots. When the robots have large motion disturbance, such as delay or deviation from the planned path, additional techniques (such as [33]) will be needed to ensure collision-free execution of the paths, which is another research topic and is not the focus of this work. In our experiments, the robots run in slow speed and the motion disturbance of the robots are relatively small.
Figure 5: Test settings for ROSbot experiments. (a) and (b) show the real-world setup. (c) and (d) show the rviz window corresponding to (a) and (b) respectively. Intermediate targets are highlighted using pink boxes and the goals using green.
We then compare the performance of DMS* (which minimizes the makespan) and MS* (which minimizes the sum of arrival times) on the mobile robots. Both planners were tested on the same map with two intermediate targets as shown in Fig. 6. We observe that DMS* tends to distribute the targets evenly to the agents and plan paths to minimize the makespan, while MS* finds a solution wherein one robot visits all targets, and the other robot directly goes to its assigned destination. MS* seeks to minimize the sum of arrival times, which can lead to a large makespan.
@@ V-C Experiments with Mobile Robots
VI Conclusion and Future Work
This paper investigates a min-max variant of Multi-Agent Combinatorial Path Finding problem and develops DMS* algorithm to solve this problem. We analyze the properties of DMS* and test DMS* with up to 20 agents and 80 targets. We also conduct robot experiments to showcase the usage of DMS* on differential-drive robots.
For future work, one can investigate how to incorporate the existing conflict resolution techniques for MAPF into the DMS* to further improve its runtime efficiency. We also note from our robot experiments that, the uncertainty and disturbance in robot motion may affect the execution of the planned path, and one can develop fast online replanning version of DMS* to address the uncertainty and disturbance.
Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. 2120219 and 2120529. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.

### v2

@@ IV-B DMS* Algorithm
After generating the successor labels $L_{succ}$ of $l$, for each $l^{\prime}\in L_{succ}$, DMS* checks for conflicts between agents during the transition from $v(l)$ to $v(l^{\prime})$, and store the subset of agents in conflict in the conflict set $I_{C}(l^{\prime})$. DMS* then invokes BackProp (Alg. 3) to back propagate $I_{C}(l^{\prime})$ to its ancestor labels recursively so that the conflict set of these ancestor labels are modified, and labels with modified conflict set are re-added to OPEN and will be re-expanded. DMS* maintains a back_set($l$) for each $l$, which is a set of pointers pointing to the predecessor labels to which the back propagation should be conducted. Intuitively, similarly to [7, 5], the conflict set of labels are dynamically enlarged during planning when agents are detected in conflict. The conflict sets of labels determine the sub-graph within the joint graph $\mathcal{G^{W}}$ that can be reached by DMS*, and DMS* always attempts to limit the search within a sub-graph of $\mathcal{G^{W}}$ as small as possible.
@@
Afterwards, if $I_{C}(l^{\prime})\neq\emptyset$, $l^{\prime}$ leads to a conflict and is discarded. Otherwise, $l^{\prime}$ is checked for pruning by using dominance (Def. 3) against any existing labels in $\mathcal{F}(v(l^{\prime}))$. If $l^{\prime}$ is pruned, any future joint path from $l^{\prime}$ can be cut and paste to $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates $l^{\prime}$ without worsening the cost to reach the goals. Furthermore, for each $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates or is equal to $l^{\prime}$, DomBackProp(Alg. 4) is invoked so that the conflict set $I_{C}(l^{\prime})$ is back propagated to $l$, and $l$ is added to the back_set of $l^{\prime\prime}$. By doing so, DMS* is able to keep updating the conflict set of the predecessor labels of $l^{\prime}$ after $l^{\prime}$ is pruned. This ensures the predecessor labels of $l^{\prime}$ will also be re-expanded after $l^{\prime}$ is pruned. If $l^{\prime}$ is not pruned, SimpleHeu is invoked for $l^{\prime}$ to compute a heuristic, which can be implemented by first copying $\vec{h}(l)$, where $l$ is the parent of $l^{\prime}$, and then reduce each component of the copied vector by one except for the components that are already zero. This heuristic is an underestimate of the cost-to-go since all agents can move at most one step closer to their goals in each expansion. Then, $f_{temp}(l^{\prime})$ is computed and $l^{\prime}$ is added to OPEN with $f_{temp}(l^{\prime})$ as its priority. Other related data structure including $\mathcal{F}(v(l^{\prime}))$, back_set, $parent$ are also updated correspondingly, and the iteration ends.
@@ IV-C Deferred Target Sequencing
DMS* introduces two techniques to defer the target sequencing until needed. As aforementioned, the first one uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic when a label is generated, and invokes TargetSeq only when that label is popped from OPEN for expansion.
@@
We now focus on the second technique in DMS*. In the previous MS* [5], every time when the search encounters a new label $l$ that is off-policy, inside TargetSeq, SolveMHPP is invoked for $l$ to find a joint sequence and policy from $l$, which burdens the overall computation, especially when a lot of new labels are generated due to the agent-agent conflict. Different from MS*, DMS* seeks to defer the call of SolveMHPP inside TargetSeq. For a label $l^{\prime}$, DMS* attempts to avoid calling SolveMHPP for $l^{\prime}$ by re-using the joint sequence $\gamma(l)$ of its parent label $l$ ($l$ is the parent of $l^{\prime}$). As presented in Alg. 2, on Lines 3-4, DMS* first attemps to build a policy from $l^{\prime}$ by following the joint sequence of its parent label $\gamma(l)$ and computes the corresponding cost-to-go $\vec{h}(l^{\prime})$. Agents $i\notin I_{C}(l)$ are still along their individual paths as in the previously computed policy $\phi_{\gamma(l)}$. Agents $i\in I_{C}(l)$ consider all actions as in Equation (1) and may deviate from the individual paths specified by the previously computed policy $\phi_{\gamma(l)}$. Therefore, DMS* needs to go through a check for these agents $i\in I_{C}(l)$: DMS* first computes the $f$-vector by summing up $\vec{g}(l^{\prime})$ and $\vec{h}(l^{\prime})$. Then, if $f^{i}(l^{\prime})$ of some agent $i\in I_{C}(l)$ is no larger than $f_{max}(l)$, DMS* can avoid calling SolveMHPP, since letting the agents follow $\gamma(l)$ in the future will not worsen the objective value, the makespan. Otherwise, there exists an agent $i\in I_{C}(l)$, whose corresponding $f^{i}(l^{\prime})$ is greater than $f_{max}(l)$, and in this case, DMS* cannot avoid calling SolveMHPP since there may exists another joint sequence from $l^{\prime}$, which leads to a solution joint path with better (smaller) objective value.
V Analysis
V-A Completeness
@@
This section discusses the properties of DMS*. An algorithm is complete for a problem if it finds a solution for any solvable instance, and reports failure in finite time for unsolvable instances.
Theorem 1
If SolveMHPP is complete, then DMS* is complete for MCPF-max.
Proof:
@@
Let $\mathcal{A}=\{0,1\}^{M}$ denote the set of all possible binary vectors and let $\mathcal{S}=\mathcal{G^{W}}\times\mathcal{A}$ denote the Cartesian of the joint graph $\mathcal{G^{W}}$ and $\mathcal{A}$. $\mathcal{S}$ is a finite space. DMS* uses dominance pruning and does not expand the same label twice during the search (the same label will be pruned by dominance). The policies computed by DMS* defines a sub-graph $\mathcal{G}^{W}_{sub}$ of the joint graph $\mathcal{G^{W}}$. DMS* first expands labels in this sub-graph $\mathcal{G}^{W}_{sub}$. If no conflict is detected when following the policy, then the resulting joint path is conflict-free and solves the problem. If there is a conflict, DMS* may re-expand a label with a larger conflict set. The conflict set of a label can be enlarged for at most $(N-1)$ times. Therefore, there is only a finite number of re-expansion of labels. As a result, for a unsolvable instance, DMS* enumerates all possible joint paths that starts from $(v_{o},\vec{0}^{M})$ to any reachable $(v,a)\in\mathcal{S}$ and terminates in finite time. For a solvable instance, DMS* conducts systematic search in $\mathcal{S}$ and returns a solution in finite time. ∎
The above analysis requires that the procedure SolveMHPP is complete for MHPP. Otherwise, if SolveMHPP never terminates for certain MHPP instances, DMS* will not terminate in finite time when calling SolveMHPP with those MHPP instances.
V-B Solution Optimality
@@
Assumption 2
Line 5 in Alg. 1 never returned true during the search of DMS*.
Theorem 2
@@
We now explain if Assumption 2 is violated, why DMS* loses the optimality guarantee and how to fix it. In MCPF-max, all agents are “coupled” in the space of binary vectors $\mathcal{A}$ in a sense that a target visited by one agent does not need to be visited by any other agents. As as result, when one agent changes its target sequence by visiting some target that is previously assigned to another agent, all agents may need to be re-planned in order to ensure solution optimality. It means that, when Line 5 in Alg. 1 returns true for label $l^{\prime}$, procedure SolveMHPP needs to be called for $l^{\prime}$ and SolveMHPP may return a new joint sequence $\gamma^{\prime}$. In this new sequence $\gamma^{\prime}$, it is possible that agents within $I_{C}(l)$ visits targets that are previously assigned to agents that are not in $I_{C}(l)$. However, DMS* does not let agents outside $I_{C}(l)$ to take all possible actions as defined in Equation (1). As a result, an optimal solution joint path may lie outside $\mathcal{G}^{W}_{sub}$ and DMS* thus loses the solution optimality guarantee when Assumption 2 is violated.
@@
V-C Solution Bounded Sub-Optimality
@@
Proof:
Bounded sub-optimal joint sequences lead to inflated heuristic values for labels in DMS*. As DMS* select labels based on $f$-values as A* does, inflated heuristic values lead to bounded sub-optimal solutions for the problem [22]. ∎
@@
Theorem 4
For a solvable instance, if Line 23 in Alg. 1 is replaced with $\vec{f}\leftarrow\vec{g}+(1+w)\vec{h}$ and Assumption 1 and 2 hold, then DMS* returns a $w$-bounded solution joint path.
VI Experimental Results
We implement DMS* in Python, with Google OR-Tool as the mHPP solver. Limited by our knowledge on Google OR-Tool, we only consider the following type of assignment constraints, where any agent can visit all targets and goals. We set up a 60-second runtime limit for each instance, where each instance contains the starts, targets and goals in a grid map. All tests run on a MacBook Pro with a Apple M2 Pro CPU and 16GM RAM. It is known from [7], M*-based algorithms with inflated heuristics can often scale to more agents, and we set $w=1.1$ in our tests. We set the number of targets $M=20,40,60,80$. We compare DMS*, which includes the two proposed technique to defer target sequencing, against two baselines. The first one is MS*, which does not have these two techniques. In other words, MS* here is a naive adaption of the existing MS* [5] algorithm to solve MCPF-max, by using the aforementioned Google OR-Tool to solve min-max mHPP for target sequencing. The second baseline is a iterative greedy approach which assigns one unvisited target to an agent to minimize the makespan of all agents in the current iteration while planning collision free paths.
Figure 3: The success rates and runtime of DMS* and MS* (baseline) with varying number of agents and targets in a random 32x32 map. DMS* has higher success rates and less runtime on average than MS*.
@@ VI Experimental Results
VI-A Simulation Results
VI-A1 Varying Number of Agents
We first fix the map to Random 32x32, and vary the number of agents $N=5,10,15,20$. We measure the success rates, the average runtime to solution and the average runtime for target sequencing per instance. The averages are taken over all instances, including both succeeded instances and instances where the algorithm times out. As shown in Fig. 3, DMS* achieves higher success rates and lower runtime than the baseline MS*. As $M$ increases from $20$ to $80$, both algorithms require more runtime for target sequencing. As $N$ increases from $5$ to $20$, agents have higher density and are more likely to run into conflict with each other. As a result, both algorithms time out for more instances. Additionally, MS* spends almost all of its runtime in target sequencing, which indicates the computational burden caused by the frequent call of mHPP solver. In contrast, for DMS*, when $N=5,10$, DMS* spends most of the runtime in target sequencing, while as $N$ increases to $15,20$, DMS* spends more runtime in path planning.
VI-A2 Different Maps
We then fixed $N=10$ and test in maps of different sizes (Fig. 4). DMS* outperforms MS* in success rates due to the alleviated computational burden for target sequencing. Larger maps do not lead to lower success rates since larger maps can reduce the density of the agents and make the agents less likely to run into conflicts with each other. The Room 32x32 map is more challenging than the other two maps since there are many narrow corridors which often lead to conflicts between the agents.
$M$ min. CR median CR max.CR
$20$ 0.47 0.71 0.94
$40$ 0.46 0.63 0.86
$60$ 0.53 0.66 0.76
$80$ 0.56 0.72 0.87
TABLE I: Cost ratios (CR) of DMS* over a greedy baseline. Solution costs of DMS* are up to 50% cheaper than the greedy.
VI-A3 Solution Quality
We use the random 32x32 map and fixed the number of agents to $N=5$ where DMS* achieves 100% success rate. We compare the solution cost ratio of DMS* over the greedy baseline. As shown in Table I, the solution of DMS* is up to around 50% cheaper than the solution of this greedy baseline.
$M$ 20 40 60 80
MS* 72.2 67.6 75.0 71.8
DMS* 77.2 84.4 93.4 99.9
TABLE II: The average number of expansion of MS* and DMS* among the instances where both algorithms succeed.
VI-A4 Number of Expansions
Table II shows the average number of label expansion in both MS* and DMS* among the instances in the Random 32x32 map which are successfully solved by both algorithms within the runtime limit. Due to the deferred sequencing, DMS* tends to search in a less informed manner and needs more expansion.
VI-B Experiments with Mobile Robots
Figure 5: Test settings for ROSbot experiments. (a) and (b) show the setup. (c) and (d) show the rviz window corresponding to (a) and (b) respectively. Targets are highlighted using pink boxes and the goals using green.
@@ VI-B Experiments with Mobile Robots
We also performed experiments with Husarion ROSbot 2, which are differential drive robots equipped with an ASUS Tinker Board and use a 2D lidar for localization. We used two such robots to perform tests in two different environments, as shown in Fig. 5 (a) and (b) having 2 and 3 targets, respectively. This test verifies that the paths planned by DMS* are executable on robots. When the robots have large motion disturbance, such as delay or deviation from the planned path, additional techniques (such as [23]) will be needed to ensure collision-free execution of the paths, which is another research topic and is not the focus of this work. In our experiments, the robots run in slow speed and the motion disturbance of the robots are relatively small.
We then compare the performance of DMS* and MS* on the robots. Both planners were tested on the same map with two intermediate targets as shown in Fig. 6. We observe that DMS* tends to assign the targets evenly to the agents and plan paths to minimize the makespan, while MS* finds a solution wherein one robot visits all targets, and the other robot directly goes to its goal. MS* seeks to minimize the sum of arrival times, which can lead to a large makespan.
VII Conclusion and Future Work
This paper investigates a min-max variant of Multi-Agent Combinatorial Path Finding problem and develops DMS* algorithm to solve this problem. We test DMS* with up to 20 agents and 80 targets and conduct simple robot experiments to showcase the usage of DMS*.
For future work, one can investigate extending DMS* to the case where edges have non-unitary traversal times [19, 20], or simultaneously optimizing both the maximum and the sum of individual arrival times [24]. We note from our experiments the disturbance in robot motion may affect the execution of the planned path, and one can develop fast online replanning version of DMS* to handle the disturbance.